OpenStudy (anonymous):

Can someone help me solve this equation? -4(-x-2) +5 = -2(2x+7)

OpenStudy (anonymous):

4x + 8 + 5 = -4x - 14 4x + 13 = -4x - 14 8x = -27 x= -27/8

OpenStudy (anonymous):

was gonna say that